# Graph-Enhanced EEG Foundation Model

### v1

@@ Introduction
Nowadays, the success of large language models (LLMs) has driven the development and widespread adoption of foundation models. These models leverage extensive amounts of unlabeled data during pre-training, achieving strong performance across a wide range of downstream tasks with minimal reliance on labeled data. Furthermore, foundation models eliminate the need for training from scratch for each task, significantly reducing computational time and cost. A notable characteristic of foundation models is their “emergent ability,” where larger architectures trained on vast datasets exhibit enhanced generalization capabilities. As a result, foundation models have gained traction across various domains, including language (Touvron et al. 2023a, b), image (Wang et al. 2023; Yuan et al. 2021), and are now being explored for EEG analysis.
@@
Several foundation models for EEG analysis, such as BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021) and Brant (Zhang et al. 2024), have been proposed, primarily treating EEG as a collection of time-series data for individual channels (electrodes). These models often rely on advanced sequence modeling techniques, including Transformers and autoencoders, to process temporal information. However, EEG signals inherently involve complex interactions between channels, making it essential to incorporate inter-channel relationships into the modeling process. For instance, abnormalities in these interactions can serve as indicators of neurological disorders, highlighting the potential diagnostic value of capturing such relationships. Despite this, existing research has not yet adequately addressed the importance of modeling contextual relationships across EEG channels.
@@
To realize our proposed approach, we build upon BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021), a masked autoencoder-based model recognized for its strong performance and adaptability for customization. In our architecture, we integrate GNNs to enhance the model’s ability to capture inter-channel relationships by representing EEG data as a graph structure. Specifically, each EEG channel is treated as a node, the signals from the corresponding channels serve as node features, and the connections between channels are modeled as edges.
@@
To address the research questions, we conducted experiments using three GNN architectures: Graph Convolutional Networks (GCN) (Kipf and Welling 2016), Graph Attention Networks (GAT) (Veličković et al. 2017), and GraphSAGE (Hamilton, Ying, and Leskovec 2017). The models were pre-trained on a large-scale dataset and evaluated on three downstream tasks to assess their performance. To account for the variability in EEG sequence lengths, we examined two adjustment strategies. Furthermore, we tested of two configurations of BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021), which have demonstrated high accuracy in prior studies, as base models integrated with GNNs to develop a robust foundation model for EEG analysis. The results show that our proposed model, particularly when employing the GCN architecture with optimized configurations, consistently outperformed baseline methods across all tasks.
@@
Background and Related Work
Foundation Models
Foundation models are large-scale pre-trained models designed to serve as general-purpose frameworks across a wide range of downstream tasks. These models are typically trained on extensive datasets using self-supervised learning, enabling the use of large amounts of unlabeled data. By leveraging their learned representations, foundation models can be adapted to specific tasks with minimal fine-tuning, demonstrating strong performance and versatility. This approach reduces the time and cost associated with task-specific model development. Foundation models have shown significant potential across various domains, including natural language processing (NLP) (Touvron et al. 2023a, b) and computer vision (CV) (Wang et al. 2023; Yuan et al. 2021).
EEG Foundation Models
Foundation models have recently been introduced in the EEG field to address key challenges, such as the difficulty of large-scale data labeling. Among these, BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021) is a notable example, leveraging techniques inspired by BERT (Devlin 2018) and wav2vec (Baevski et al. 2020). However, to the best of our knowledge, existing EEG foundation models focus exclusively on learning time-series information, while neglecting inter-channel relationships, which are also critical for capturing the underlying dynamics of EEG signals.
EEG Models with Graphs
Several EEG models have been developed to capture inter-channel relationships, although these are not foundation models but are instead trained from scratch for specific tasks. These models represent EEG signals as graphs, allowing the study of network properties such as channel connectivity. To process these graph-structured data, they employ Graph Neural Networks (GNNs), which are specifically designed for such applications (Klepl, Wu, and He 2024). One notable example is EEG-GCNN (Wagh and Varatharajah 2020), which utilizes a Graph Convolutional Network (GCN)  (Kipf and Welling 2016) architecture with edge weights calculated based on spatial distances. This model learns inter-channel relationships effectively by leveraging these edge weights.
Graph-Enhanced Models in Other Domains
@@
Pre-trained language models like BERT (Devlin 2018) have demonstrated exceptional performance across various NLP tasks. These models are typically trained using self-supervised learning techniques, such as masked language modeling, during pre-training. Building upon such models, LinkBERT (Yasunaga, Leskovec, and Liang 2022) extends the input context beyond a single document by modeling links between documents.
This approach aligns with our interest in treating data as a graph structure, where node features are processed with powerful architectures while also learning the connectivity between nodes. LinkBERT achieves this by introducing an additional pretext task during pre-training: predicting whether two document segments are linked (in the same context) or not.
However, this method is not directly applicable to EEG data, as EEG signals exhibit less variation compared to textual data, making it difficult to design a similar pretext task for identifying relationships between signals. Instead, we propose the use of GNNs, which can effectively model connectivity based on spatial information and are better suited to learning inter-channel relationships in EEG.
Methods
We propose a foundation model that learns both the inter-channel relationships and the time-series dynamics of EEG signals. Since existing EEG foundation models primarily focus on learning time-series information, we extend these models by integrating inter-channel relationship learning. Among the state-of-the-art EEG foundation models, BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021) is notable for its strong performance and adaptability for customization. Therefore, we adopt BENDR as the base foundation model for our proposed approach. In this section, we first provide a brief overview of BENDR, followed by the description of our proposed method.
Figure 1: Comparison between two model architectures during pre-training and fine tuning. Figures (a), (c) and (e) represent BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021), while (b), (d) and (f) represent our proposed model.
BENDR
@@ Methods
During pre-training, raw multi-channel EEG signals are passed through six convolutional layers (referred to as the “BENDR Encoder”) to generate convolved features. The procedure is inspired by wav2vec (Baevski et al. 2020). A portion of the convolved features is then masked, and a Transformer encoder reconstructs the masked features using the unmasked features as context. The model is optimized by calculating the contrastive loss between the reconstructed features and the original convolved features before masking.
For downstream tasks, several model configurations have been proposed. Among these, we focus on two configurations (“BENDR” and “Linear”) that achieved the highest performance according to (Kostas, Aroca-Ouellette, and Rudzicz 2021). In both configurations, the process up to encoding by the “BENDR Encoder” is identical to that of the pre-training phase. However, the subsequent steps differ. In the “BENDR” configuration, the convolved features are passed to the pre-trained Transformer encoder, followed by a linear layer for classification. In contrast, in the “Linear” configuration, the pre-trained Transformer encoder is omitted entirely, and the convolved features are aggregated and passed to a linear classification layer.
Our Proposal
@@
  • Following EEG-GCNN (Wagh and Varatharajah 2020), the graph is fully connected, and each edge $(u,v)\in E$ has a weight. This setup enables the GNN to capture all pairwise connectivity between channels. The edge weights are represented as a matrix $W\in\mathbf{R}^{C\times C}$.
@@
Experiments
@@
Dataset
Pre-training Dataset
Following BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021), we utilized the Temple University Hospital EEG Corpus (TUEG) (Obeid and Picone 2016) for pre-training. TUEG provides a diverse range of subjects and includes recordings across multiple sessions over extended time periods, making it an ideal dataset for pre-training foundation models. For our study, we specifically focused on version 2 of this dataset, which consists of clinical recordings from over 10,000 individuals. To accelerate experimentation and evaluation during development, we downsampled the dataset to one-tenth of its original size.
Downstream Datasets
@@
(Goldberger et al. 2000; Schalk et al. 2004) This task involves predicting whether the participant is imagining the movement of the right (positive) or left (negative) hand.
@@
(Goldberger et al. 2000; Citi, Poli, and Cinel 2010) This task involves predicting whether the participant focused on a flashing target letter (positive) or a non-target letter (negative).
@@
(Margaux et al. 2012) This task involves predicting whether the participant’s attempt to input a character using a P300 speller was recognized correctly (positive) or incorrectly (negative).
All three tasks are binary classification problems, which were previously used in BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021). These tasks involve EEG signals recorded from a sufficient number of channels in the 10/20 channel scheme (Jurcak, Tsuzuki, and Dan 2007), making them suitable for constructing graphs for GNN-based approaches. Detailed information about these datasets is provided in Table 1.
The P300 and ERN datasets exhibit imbalanced class distributions, while the MMI dataset is balanced. To address class imbalance during fine tuning, we followed the methodology presented in BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021) and applied the following steps:
@@
MMI (Goldberger et al. 2000; Schalk et al. 2004) 160 6 64 105 5
ERN (Goldberger et al. 2000; Citi, Poli, and Cinel 2010) 200 2 56 26(10) 4
P300 (Margaux et al. 2012) 2,048 2 64 9 9
Table 1: Downstream dataset battery and number of cross-validation folds used, following BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021).
Preprocessing
We applied the following preprocessing steps to both the pre-training and downstream datasets, following the methodology presented in BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021):
@@
  • We utilized 19 EEG channels from the 10/20 channel scheme (Jurcak, Tsuzuki, and Dan 2007) and ignored all other channels.
  • For pre-training, 60-second sequences were extracted from the pre-training dataset for use in general training, while 20-second sequences were specifically used for the P300-related training, as described in BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021). For downstream tasks, the entire length of the sequences was used.
Setup
GNN architectures
We evaluated the performance of our proposed model by individually incorporating three standard GNN architectures: Graph Convolutional Networks (GCN) (Kipf and Welling 2016), Graph Attention Netowrks (GAT) (Veličković et al. 2017) and GraphSAGE (Hamilton, Ying, and Leskovec 2017). Each architecture was used consistently throughout a single set of pre-training and downstream tasks. All GNN implementations were based on PyTorch 2.3.1 and PyTorch Geometric 2.5.3.
@@
Sequence Length Adjusters
@@
Results
The results are presented in Tables 2 and 3. The baseline corresponds to the original BENDR (Kostas, Aroca-Ouellette, and Rudzicz 2021). Note that the evaluation metric for MMI is Accuracy, while AUROC is used for P300 and ERN. The abbreviation “(w/ e.)” in the tables indicates that the GNN utilized edge weights. The following discussion addresses key points related to our research questions.
Comparison of GNN Architectures (RQ1)
Tables 2 and 3 indicate that our proposed approach performed better when using a linear layer for the sequence length adjustment compared to padding. Therefore, to analyze the performance variations introduced by different GNN architectures, the following discussion focuses on the results obtained with a linear layer.
As shown in Table 2, among all the GNN architectures tested, only “GCN with edge weights” consistently outperformed the baseline across all three downstream tasks. Thus, “GCN with edge weights” emerges as the most suitable architecture for incorporation into foundation models for EEG analysis. The next best-performing architecture is “GraphSAGE”, which exceeded the baseline in two out of three tasks, making it another promising candidate. Investigating the underlying factors contributing to this behavior remains an open question and is expected to be addressed in future work.
@@ Results
Table 2: The results for all downstream tasks and GNN architectures using a linear layer as the sequence length adjuster. Entries from our proposed models that exceed the baseline are underlined.
Differences of the Effect of GNNs Arising from Downstream Tasks (RQ2)
@@
As shown in Table 2, on MMI, the baseline achieved a score of $0.646$, while our model achieved $0.849$, representing a $31.4\%$ improvement. On P300, the baseline achieved $0.568$, with our model improving performance by $8.53\%$. On ERN, the baseline achieved $0.522$, and our model showed a $3.11\%$ improvement. These results indicate that the higher the baseline performance on a task, the greater the relative improvement achieved by our model. Further analysis could investigate the relationship between model performance and task-specific characteristics, particularly the physiological features associated with each task, to gain deeper insights.
Comparison of Sequence Length Adjusters (RQ3)
Tables 2, 3 show that all the models using a linear layer as the sequence length adjuster consistently outperformed those using padding. Consequently, more models outperformed the baseline when using a linear layer compared to padding. This indicates that adding a linear layer before GNNs is a more effective method for sequence length adjustment.

### v2

@@ I Introduction
Nowadays, the success of large language models (LLMs) has driven the development and widespread adoption of foundation models. These models leverage extensive amounts of unlabeled data during pre-training, achieving strong performance across a wide range of downstream tasks with minimal reliance on labeled data. Furthermore, foundation models eliminate the need for training from scratch for each task, significantly reducing computational time and cost. A notable characteristic of foundation models is their “emergent ability,” where larger architectures trained on vast datasets exhibit enhanced generalization capabilities. As a result, foundation models have gained traction across various domains, including language [1, 2], image [3, 4], and are now being explored for EEG analysis.
@@
Several foundation models for EEG analysis, such as BENDR [5] and Neuro-GPT [6], have been proposed, primarily treating EEG as a collection of time-series data for individual channels (electrodes). These models often rely on advanced sequence modeling techniques, including Transformers and autoencoders, to process temporal information. However, EEG signals inherently involve complex interactions between channels, making it essential to incorporate inter-channel relationships into the modeling process. For instance, abnormalities in these interactions can serve as indicators of neurological disorders, highlighting the potential diagnostic value of capturing such relationships. Despite this, existing research has not yet adequately addressed the importance of modeling contextual relationships across EEG channels.
@@
To realize our proposed approach, we build upon BENDR [5], a masked autoencoder-based model recognized for its strong performance and adaptability for customization. In our architecture, we integrate GNNs to enhance the model’s ability to capture inter-channel relationships by representing EEG data as a graph structure. Specifically, each EEG channel is treated as a node, the signals from the corresponding channels serve as node features, and the connections between channels are modeled as edges.
@@
To address the research questions, we conducted experiments using three GNN architectures: Graph Convolutional Networks (GCN) [7], Graph Attention Networks (GAT) [8], and GraphSAGE [9]. The models were pre-trained on a large-scale dataset and evaluated on three downstream tasks to assess their performance. To account for the variability in EEG sequence lengths, we examined two adjustment strategies. Furthermore, we tested of two configurations of BENDR [5], which have demonstrated high accuracy in prior studies, as base models integrated with GNNs to develop a robust foundation model for EEG analysis. The results show that our proposed model, particularly when employing the GCN architecture with optimized configurations, consistently outperformed baseline methods across all tasks.
@@
II Background and Related Work
II-A Foundation Models
Foundation models are large-scale pre-trained models designed to serve as general-purpose frameworks across a wide range of downstream tasks. These models are typically trained on extensive datasets using self-supervised learning, enabling the use of large amounts of unlabeled data. By leveraging their learned representations, foundation models can be adapted to specific tasks with minimal fine-tuning, demonstrating strong performance and versatility. This approach reduces the time and cost associated with task-specific model development. Foundation models have shown significant potential across various domains, including natural language processing (NLP) [1, 2] and computer vision (CV) [3, 4].
II-B EEG Foundation Models
Foundation models have recently been introduced in the EEG field to address key challenges, such as the difficulty of large-scale data labeling. Among these, BENDR [5] is a notable example, leveraging techniques inspired by BERT [10] and wav2vec [11]. First, raw multi-channel EEG signals are passed through six convolutional layers (referred to as the “BENDR Encoder”) to generate convolved features. A portion of the convolved features is then masked, and a Transformer encoder reconstructs the masked features using the unmasked features as context. The model is optimized by calculating the contrastive loss between the reconstructed features and the original convolved features before masking.
MAEEG [12] is an improved model that builds upon the BENDR [5] method, achieving enhanced performance. Specifically, a linear layer and a convolutional layer are added after the Transformer encoder to directly reconstruct the EEG signal. And then instead of computing the contrastive loss between the original features and the reconstructed features, MAEEG calculates the reconstruction loss between the input EEG signal and the reconstructed EEG signal.
Neuro-GPT [6] is another model that has achieved higher performance than BENDR [5]. This model is based on the Generative Pre-trained Transformer (GPT) [13] method and uses a decoder-only Transformer Encoder for the encoded features by a new EEG Encoder, which consists of a convolutional module and a Transformer Encoder.
However, to the best of our knowledge, existing EEG foundation models focus exclusively on learning time-series information, while neglecting inter-channel relationships, which are also critical for capturing the underlying dynamics of EEG signals.
II-C EEG Models with Graphs
Several EEG models have been developed to capture inter-channel relationships, although these are not foundation models but are instead trained from scratch for specific tasks. These models represent EEG signals as graphs, allowing the study of network properties such as channel connectivity. To process these graph-structured data, they employ Graph Neural Networks (GNNs), which are specifically designed for such applications [14]. One notable example is EEG-GCNN [15], which utilizes a Graph Convolutional Network (GCN)  [7] architecture with edge weights calculated based on spatial distances. This model learns inter-channel relationships effectively by leveraging these edge weights.
II-D Graph-Enhanced Models in Other Domains
@@
One example is the pre-training language models (LM) with knowledge graphs (KGs), such as K-BERT [16], which incorporates knowledge graphs into the pre-training process to enhance the model’s understanding of entities and relationships. Some research has proposed the combination of LMs and GNNs to jointly learn on texts and KGs [17, 18, 19]. We can apply this approach of integrating GNNs with LMs to EEG data, where the GNNs can capture the relationships between EEG channels, while the LMs can be replaced with Transformer-based sequence models to learn the temporal dynamics of the signals.
III Methods
We propose a foundation model that learns both the inter-channel relationships and the time-series dynamics of EEG signals. Since existing EEG foundation models primarily focus on learning time-series information, we extend these models by integrating inter-channel relationship learning. Among the state-of-the-art EEG foundation models, BENDR [5] is notable for its strong performance and adaptability for customization. Therefore, we adopt BENDR as the base foundation model for our proposed approach. In this section, we first provide a brief overview of BENDR, followed by the description of our proposed method.
Figure 1: Comparison between two model architectures during pre-training and fine tuning. Figures (a), (c) and (e) represent BENDR [5], while (b), (d) and (f) represent our proposed model.
III-A BENDR
@@
During pre-training, raw multi-channel EEG signals are passed through six convolutional layers (referred to as the “BENDR Encoder”) to generate convolved features. The procedure is inspired by wav2vec [11]. A portion of the convolved features is then masked, and a Transformer encoder reconstructs the masked features using the unmasked features as context. The model is optimized by calculating the contrastive loss between the reconstructed features and the original convolved features before masking.
For downstream tasks, several model configurations have been proposed. Among these, we focus on two configurations (“BENDR” and “Linear”) that achieved the highest performance according to [5]. In both configurations, the process up to encoding by the “BENDR Encoder” is identical to that of the pre-training phase. However, the subsequent steps differ. In the “BENDR” configuration, the convolved features are passed to the pre-trained Transformer encoder, followed by a linear layer for classification. In contrast, in the “Linear” configuration, the pre-trained Transformer encoder is omitted entirely, and the convolved features are aggregated and passed to a linear classification layer.
III-B Our Proposal
@@
  • Following EEG-GCNN [15], the graph is fully connected, and each edge $(u,v)\in E$ has a weight. This setup enables the GNN to capture all pairwise connectivity between channels. The edge weights are represented as a matrix $W\in\mathbf{R}^{C\times C}$.
@@
III-C Alternative Architectural Variants
Our proposed method aims to learn both temporal and spatial information in EEG signals as a foundation model. To achieve this, we leverage the strengths of existing EEG foundation models in learning temporal information and add a module to learn spatial relationships between channels.
There are three possible ways to combine these modules:
  1. 1.
    Learning spatial relationships first and then temporal information (This is the method we adopted, as described above).
  2. 2.
    Learning temporal information first and then spatial relationships.
  3. 3.
    Learning both in parallel and aggregating the results.
However, approach 2 is not suitable for our purpose because it would require placing the GNN after the BENDR Encoder and Transformer Encoder, which would integrate all channel information and make it difficult to learn spatial relationships.
Approach 3 is thought to be feasible, but we chose to focus on approach 1 in this study because it allows us to reuse the BENDR Encoder and pretext task with minimal modifications. Approach 3 is worth exploring in future research.
In this study, we adopt approach 1, where we first learn spatial relationships using a GNN and then learn temporal information using the BENDR Encoder and Transformer Encoder. This approach enables us to effectively learn both spatial and temporal information in EEG signals.
III-D Incorporating Graphs Into Other EEG Foundation Models
In addition to BENDR [5], there also exist other EEG foundation models. Although we did not conduct experiments on them, we believe that a similar approach can be applied to them as well. Here we provide a brief overview of how our proposed method can be applied to some other EEG foundation models, such as MAEEG [12] and Neuro-GPT [6].
Both models first use a Convolutional Encoder to process the input EEG signals and then use a Transformer Encoder to capture temporal information. We consider a simple yet effective method, inserting a GNN before the Convolutional Encoder to learn the relationships between channels, will be effective for these models. By incorporating graph structures into these EEG foundation models, we can potentially improve their performance and better capture the complex relationships between channels in EEG data.
IV Experiments
@@
IV-A Dataset
IV-A1 Pre-training Dataset
Following BENDR [5], we utilized the Temple University Hospital EEG Corpus (TUEG) [20] for pre-training. TUEG provides a diverse range of subjects and includes recordings across multiple sessions over extended time periods, making it an ideal dataset for pre-training foundation models. For our study, we specifically focused on version 2 of this dataset, which consists of clinical recordings from over 10,000 individuals. To accelerate experimentation and evaluation during development, we downsampled the dataset to one-tenth of its original size.
IV-A2 Downstream Datasets
@@
[21, 22] This task involves predicting whether the participant is imagining the movement of the right (positive) or left (negative) hand.
@@
[21, 23] This task involves predicting whether the participant focused on a flashing target letter (positive) or a non-target letter (negative).
@@
[24] This task involves predicting whether the participant’s attempt to input a character using a P300 speller was recognized correctly (positive) or incorrectly (negative).
All three tasks are binary classification problems, which were previously used in BENDR [5]. These tasks involve EEG signals recorded from a sufficient number of channels in the 10/20 channel scheme [25], making them suitable for constructing graphs for GNN-based approaches. Detailed information about these datasets is provided in Table I.
The P300 and ERN datasets exhibit imbalanced class distributions, while the MMI dataset is balanced. To address class imbalance during fine tuning, we followed the methodology presented in BENDR [5] and applied the following steps:
@@
MMI [21, 22] 160 6 64 105 5
ERN [21, 23] 200 2 56 26(10) 4
P300 [24] 2,048 2 64 9 9
TABLE I: Downstream dataset battery and number of cross-validation folds used, following BENDR [5].
IV-A3 Preprocessing
We applied the following preprocessing steps to both the pre-training and downstream datasets, following the methodology presented in BENDR [5]:
@@ IV-A3 Preprocessing
  • We utilized 19 EEG channels from the 10/20 channel scheme [25] and ignored all other channels.
  • For pre-training, 60-second sequences were extracted from the pre-training dataset for use in general training, while 20-second sequences were specifically used for the P300-related training, as described in BENDR [5]. For downstream tasks, the entire length of the sequences was used.
IV-B Setup
IV-B1 GNN architectures
We evaluated the performance of our proposed model by individually incorporating three standard GNN architectures: Graph Convolutional Networks (GCN) [7], Graph Attention Netowrks (GAT) [8] and GraphSAGE [9]. Each architecture was used consistently throughout a single set of pre-training and downstream tasks. All GNN implementations were based on PyTorch 2.3.1 and PyTorch Geometric 2.5.3.
@@ IV-B1 GNN architectures
IV-B2 Sequence Length Adjusters
@@
IV-C Results
The results are presented in Tables II and III. The baseline corresponds to the original BENDR [5]. Note that the evaluation metric for MMI is Accuracy, while AUROC is used for P300 and ERN. Those with the statement “(with edge weights)” in the tables indicates that the GNN utilized edge weights. For the experiments with padding as the sequence length adjuster (shown in Table III), GCN and GAT without edge weights were excluded from these experiments due to their poor performance with the linear layer (see Table II). The following discussion addresses key points related to our research questions.
IV-C1 Comparison of GNN Architectures (RQ1)
Tables II and III indicate that our proposed approach performed better when using a linear layer for the sequence length adjustment compared to padding. Therefore, to analyze the performance variations introduced by different GNN architectures, the following discussion focuses on the results obtained with a linear layer.
As shown in Table II, among all the GNN architectures tested, only “GCN with edge weights” consistently outperformed the baseline across all three downstream tasks. Thus, “GCN with edge weights” emerges as the most suitable architecture for incorporation into foundation models for EEG analysis. The next best-performing architecture is “GraphSAGE”, which exceeded the baseline in two out of three tasks, making it another promising candidate. Investigating the underlying factors contributing to this behavior remains an open question and is expected to be addressed in future work.
@@ IV-C1 Comparison of GNN Architectures (RQ1)
TABLE II: The results for all downstream tasks and GNN architectures using a linear layer as the sequence length adjuster.
IV-C2 Differences of the Effect of GNNs Arising from Downstream Tasks (RQ2)
@@
As shown in Table II, on MMI, the baseline achieved a score of $0.646$, while our model achieved $0.849$, representing a $31.4\%$ improvement. On P300, the baseline achieved $0.568$, with our model improving performance by $8.53\%$. On ERN, the baseline achieved $0.522$, and our model showed a $3.11\%$ improvement. These results indicate that the higher the baseline performance on a task, the greater the relative improvement achieved by our model. Further analysis could investigate the relationship between model performance and task-specific characteristics, particularly the physiological features associated with each task, to gain deeper insights.
IV-C3 Comparison of Sequence Length Adjusters (RQ3)
Tables II, III show that all the models using a linear layer as the sequence length adjuster consistently outperformed those using padding. Consequently, more models outperformed the baseline when using a linear layer compared to padding. This indicates that adding a linear layer before GNNs is a more effective method for sequence length adjustment.